# Josephson effect and self-trapping in helicoidal spin-orbit coupled Bose-Einstein condensates with optical lattices

### v1

@@ Abstract
We consider helicoidal spin-orbit coupled Bose-Einstein condensates in optical lattices and study the effects of helicoidal gauge potential on the population imbalance between two pseudo-spin states for zero and non-zero Zeeman splitting. We see that the population imbalance executes Josephson-type oscillation for zero Zeeman splitting and the frequency of oscillation increases with the decrease of strength of helicoidal gauge potential. In presence of optical lattices, the frequency further increases. However, the increment of frequency depends on wavenumber of the lattice. For non-zero Zeeman splitting, we find that the oscillation of population imbalance is not symmetric about zero imbalance of the two states. The asymmetry increases with the increase of helicoidal gauge potential and Zeeman splitting. However, it decreases with the increase of initial phase difference.
@@ I Introduction
Josephson physics describes the tunnelling phenomena that arises due to the coupling between two macroscopically phase coherent quantum states such as superconductors separated by thin insulating layer and, weakly coupled Bose-Einstein condensates(BEC) separated by tunnelling layer. The interplay between tunnelling and interaction results supercurrent due to the oscillation of population difference between these two states [1, 2]. This is the so-called Josephson oscillation. A transition from Josephson oscillation to self-trapping occurs if interaction dominates over the tunnelling and causes the population imbalance to oscillate about a non-zero value. This tunnelling oscillation can be explored in building quantum mechanical circuit to design qubit [3], SQUIDS [4] and to engineer quantum states [5].
@@
Two Bose-Einstein condensates in a trap consisting of atoms differing in hyperfine states provide a platform to realize bosonic Josephson junction of a weakly interacting system. A spin-orbit coupled (SOC) atomic BEC created by artificial gauge potential offers a new possibility to study Josephson oscillation in pseudo-spin 1/2 system. The Josephson oscillation in a weakly interacting BEC is a non-linear generalization of typical a.c. and d.c. Josephson oscillations. Thus it has received special attention and results several interesting experimental [6, 7, 8] and theoretical [9, 10, 11, 12, 13, 14, 15, 16, 17] investigations. Particularly, theoretical investigations in this direction include the studies on Josephson dynamics in two species BECs in double-well potential [10], internal Josephson oscillation between coupled solitons with time varying Raman frequency [17], effects of SOC on Josephson tunnelling between two annular condensates [18], Josephson oscillation of 2D chiral solitons [19] and Josephson-type oscillation in presence of optical lattices [20].
@@
Besides the above mentioned Josephson effects in real space, it has also been predicted theoretically [21] and, realised experimentally [22] momentum-space Josephson effect in SOC-BEC. Studies in this direction are not only limited to atomic-bosonic systems but also extended from exciton-polariton system [23] to fermionic superfluid gas [24, 25].
@@
Experimental techniques for generating various artificial gauge potentials allow to tune strength and shape of spin-orbit coupling in BEC [26, 27, 28]. For the choice of gauge potential, $A(x)=\sigma n(x)$, where $n(x)=(\cos(2\beta x),\sin(2\beta x),0)$ and $\sigma=(\sigma_{1},\sigma_{2},\sigma_{3})$, one can induce helicoidal shaped non-uniform spin-orbit coupling in BEC. This non-uniform SOC is highly tunable and can increase the effects of intrinsic non-linearity in BEC [26, 27, 28, 29, 30, 31]. A Bose-Einstein condensate with helicoidal SOC features many properties of integrable system such as stable propagation of bright solitons [32], Peregrine solitons [33], bright-dark Peregrine solitons [34], stripe solitons [35], multi-hump solitons [36], and kink solitons [37]. One of the important aspects of the soliton in helicoidal SOC-BEC is that the soliton with any peak amplitude is non-radiative and thus can propagate a large distance [32]. In a recent study it is shown that the helicoidal gauge potential can be used to manipulate modulational instability in BEC [38].
@@
Our objective in this paper is to study Josephson-type(JT) oscillation in helicoidal spin-orbit coupled quasi one-dimensional(Q1D) BECs with optical lattices. In a Q1D-BEC with traditional spin-orbit coupling (without helicoidal gauge potential) the JT oscillation has been studied for equal Rashba and Dresselhaus couplings. The helicoidal-SOC introduces inequality between these two couplings. We see that the number of atoms oscillates periodically between the two solitons and the frequency of oscillation depends sensitively on the strength of helicoidal gauge potential. Specifically, the frequency decreases with the increase of helicoidal gauge potential. In presence of optical lattices, the frequency of oscillation further increases with the decrease of lattice wavenumber [20]. Interestingly, we see that the population imbalance oscillates about its zero value if the Zeeman splitting is zero leading to JT oscillation. For a non-zero value of Zeeman splitting, the oscillation is found to result self-trapping.
@@
In section II, we present a theoretical model to describe helicoidal SOC-BEC in optical lattices for zero and non-zero Zeeman splitting. Based on a variational approach we derive equations for the center of mass, population imbalance and phase difference of the coupled condensates. Here we also examine the effects of optical lattices and Zeeman splitting on the dispersion relation. In section III, we study the dynamics of population imbalance for a particular location of center of mass in the optical lattice. Specifically, we consider zero Zeeman splitting ($\Delta$=0) and study the effects of optical lattices for different strengths of the helicoidal gauge potential. In section IV, we study the effect of non-zero Zeeman splitting on the population imbalance. Particularly, it is seen that, quantum mechanical self-trapping occurs for $\Delta\neq 0$. Finally, we devote section V, to summarize the results of our investigations.
@@ II General Formalism
A Bose-Einstein condensate can be treated theoretically by a single mean field Gross-Pitaevskii equation. In presence of helicoidal spin-orbit coupling, the BEC describes a pseudo-spin 1/2 system and thus it becomes a binary BEC. This system in quasi-one-dimensional (Q1D) geometry is modelled by the following coupled Gross-Pitaevskii equations (GPE) [32, 38].
@@
Equations (1) and (2) describe a helicoidal spin-orbit coupled BEC in linear optical lattices (LOL) if the effects of harmonic trap is negligible ($\omega_{\perp}\gg\omega_{x}$). In addition to the linear optical lattice, one can introduce periodicity in the system by periodically modulating the non-linear mean field interaction, called non-linear optical lattice(NOL), with the help of optical Feshbach resonance technique [41]. The NOLs are given by
@@
where $k_{n}$ is the wavenumber of the NOLs with their strengths $\gamma_{1}$ and $\beta_{1}$ respectively. The NOLs are found very effective in controlling the dynamics and stability of matter-wave soliton in BEC [39, 40, 41, 42, 43, 44]. In order to proceed with the variational approach we recast the Eqs.(1)and (2) in terms of action principle and obtain the following Lagrangian density.
@@
-- $\frac{1}{2}\bigg{|}\frac{d\psi_{1}}{dx}\bigg{|}^{2}-\frac{1}{2}\bigg{|}\frac{d\psi_{2}}{dx}\bigg{|}^{2}-\frac{\Delta}{2}|\psi_{1}|^{2}+\frac{\Delta}{2}|\psi_{2}|^{2}-\frac{\gamma_{0}}{2}|\psi_{1}|^{4}$
-- $\frac{{\gamma_{0}}}{2}|\psi_{2}|^{4}-{\beta_{0}}|\psi_{1}|^{2}|\psi_{2}|^{2}-V_{0}\cos(2k_{l}x)|\psi_{1}|^{2}$
-- $V_{0}\cos(2k_{l}x)|\psi_{2}|^{2}-\frac{\gamma_{1}}{2}\cos(2k_{n}x)|\psi_{1}|^{4}$
-- $\frac{\gamma_{1}}{2}\cos(2k_{n}x)|\psi_{2}|^{4}-\beta_{1}\cos(2k_{n}x)|\psi_{2}|^{2}|\psi_{1}|^{2}$
@@
A helicoidal spin-orbit coupled BEC supports freely moving bright solitons both in presence and absence of Zeeman splitting [32]. In view of this, we adopt the following Gaussian-shaped trial solutions [17].
@@
-- $\frac{1}{2\sqrt{2\pi}w}(\gamma_{0}N_{1}^{2}+\gamma_{0}N_{2}^{2}+2\beta_{0}N_{1}N_{2})$
-- $\frac{e^{\frac{-k_{n}^{2}w^{2}}{2}}}{2\sqrt{2\pi}w}(\gamma_{1}N_{1}^{2}+\gamma_{1}N_{2}^{2}+2\beta_{1}N_{1}N_{2})\cos(2k_{n}x_{0})$
@@
$+$ $\frac{Nk_{n}\sin(2k_{n}x_{0})}{2\sqrt{2\pi}w}e^{\frac{-k_{n}^{2}w^{2}}{2}}[G_{+}+G_{-}Z^{2}]$
@@
$\frac{d\varphi}{dt}$ $=$ $2\beta k_{+}-\Lambda Z+\frac{2\alpha_{s}\beta e^{-\beta^{2}w^{2}}}{\sqrt{1-Z^{2}}}\cos(\varphi)$ (11)
-- $\frac{NZ}{\sqrt{2\pi}w}e^{\frac{-k_{n}^{2}\omega^{2}}{2}}\cos(2k_{n}x_{0})G_{-}$
@@
with $\Lambda=N(\gamma_{0}-\beta_{0})/(\sqrt{2\pi})w$, $k_{\pm}=(k_{1}\pm k_{2})/2$, $G_{-}=\gamma_{1}-\beta_{1}$ and $G_{+}=\gamma_{1}+\beta_{1}$. Here, $\varphi=\phi_{1}-\phi_{2}$ and $Z=(N_{1}-N_{2})/N$, represent respectively phase difference and population imbalance between the two solitons. For a weak helicoidal gauge potential we take $k_{-}\approx\beta$. The coupled Eqs.(8)-(11) show that the rate of change of population imbalance depends directly on the initial phase difference and SOC parameters. Clearly, its dependence on the lattice and other parameters comes through $\varphi$. Therefore, it will be quite interesting to check how the population imbalance leading to JT oscillation are affected by the different parameters of the system.
II.1 Linear energy spectrum
Fundamental solitons propagate in a medium due to interplay between the non-linearity and group velocity dispersion. The propagation of soliton is further affected by the other types of dispersions present in the system. The spin-orbit coupling and optical lattices induce dispersive effect on the propagation of matter-wave solitons. With a view to understand the effect of optical lattices and spin-orbit coupling, we obtain the following energy-momentum relation using plane wave solution, $(\psi_{a},\psi_{b})=(u_{0},v_{0})\exp[ikx-i\omega(k)t]$,
Figure 1: Top panel: Variation of $\omega_{\pm}(k)$ with $k$ in absence of optical lattices with $\beta=1$, $\alpha_{s}=2$ for (i) $\Delta=0$ (left figure), (ii) $\Delta=2$ (right figure). Bottom panel: Variation of $\omega_{\pm}(k)$ with $k$ in presence of optical lattices with $\beta=1$, $\alpha_{s}=1$ and $k_{l}=2$ for (i) $\Delta=0$(left figure), (ii) $\Delta=2$ (right figure).
$\omega_{\pm}(k)=\bigg{[}\frac{1}{2}k^{2}+{\sqrt{\frac{2}{\pi}}\frac{V_{0}k}{k^{2}-4k_{l}^{2}}\bigg{]}}\pm\sqrt{\alpha_{s}^{2}k^{2}+(\beta k-\Delta/2)^{2}}.$ (12)
Here the first term represents group velocity dispersion while the second term arises due to optical lattices and the third term comes from spin-orbit coupling. The variation of $\omega_{\pm}$ with $k$ displayed in Fig.1 shows two distinct branches, namely, upper branch ($\omega_{+}$, black curve) and lower branch($\omega_{-}$, blue curve). We see that the upper branch has single minimum and lower branch has double minima with same energy for zero Zeeman splitting. In presence of optical lattices two discontinuity arises at $k=\pm 2k_{l}$ on either sides of $k=0$ [20]. For a non-zero value of Zeeman splitting the upper branch has a single minimum while the lower branch has double minima. Unlike the case of $\Delta=0$, the minima in the lower branch are non-degenerate and distinct gap arises between the upper and lower branch for $\Delta\neq 0$. The propagation of non-linear mode depends crucially on the minimum($\omega_{m}$) of the dispersion curve. More specifically, a non-linear mode (e.g, soliton) can propagate in a SOC-BEC if the chemical potential lies below the energy corresponds to $\omega_{m}$ [45].
III JT oscillation for zero Zeeman splitting
@@
Let us consider that the coupled solitons are located at a particular position in the lattice potential. For these stationary solitons, the center of mass of the coupled system does not move in time and thus $\dot{x_{c}}=0$. In this case, the coupled Eqs.(8)-(11) can be reduced as follows.
$\frac{dZ}{dt}$ $=$ $A\sin(2\varphi),$ (13)
$\frac{d\varphi}{dt}$ $=$ $-\Lambda Z-\frac{AZ}{\sqrt{1-Z^{2}}}\cos^{2}(\varphi)-BZ-\Delta.$ (14)
@@
Here, $A=2\alpha_{s}^{2}e^{-2\beta^{2}w^{2}}$ and $B=\frac{NG_{-}e^{\frac{-k_{n}^{2}w^{2}}{2}}}{\sqrt{2\pi}w}\cos(2k_{n}x_{0})$. Understandably, Eqs.(13) and (14) can describe the variation of population imbalance with a change of relative phase between the two spin states both in absence and presence of optical lattices (OLs). With a view to study population imbalance dynamics, we solve Eqs.(13) and (14) numerically for different values of initial phase differences and display the results in Fig.2. In absence of optical lattices, we see that the population imbalance oscillates with time and amplitude of the oscillation depends sensitively on the initial value of $\varphi$ (left panel). More specifically, we see that amplitude increases with the increase of $\varphi(0)$ for any value of $\beta$. This type of oscillation is often termed as Josephson-type(JT) oscillation [16, 20, 17]. However, for a fixed value of $\varphi(0)$, we see that the frequency of oscillation decreases with the increase of $\beta$ (Fig.2, right panel) keeping the amplitude same.
@@
Figure 2: Left panel: Variation of population imbalance with $t$ in absence of optical lattices for $\beta=1.6$. Here, solid black and dashed black are drawn for the initial phase differences $\varphi(0)=\pi/8$ and $\pi/16$ respectively. Right panel: It gives $Z(t)$ versus $t$ for $\beta=1.6$(solid black) and $\beta=2$ (black dashed) with $\varphi(0)=\pi/8$. In both the panels we take $\alpha_{s}=4$, $x_{0}=0.25$ and $\Delta=0$.
@@
Figure 3: Left panel: Variation of population imbalance with $t$ in presence of optical lattices for $\beta=1.6$. Here, solid black and dashed black are drawn for the initial phase differences $\varphi(0)=\pi/8$ and $\varphi(0)=\pi/16$. Right panel: It gives $Z(t)$ versus $t$ for $\beta=1.6$(solid black) and $\beta=2$ (black dotted) with $\varphi=\pi/8$. In both the panels, we take $\alpha_{s}=4$, $x_{0}=0.25$, $G_{-}=1.05$, $k_{n}=1.2$ and $\Delta=0$.
Figure 4: Left panel: Variation of JT oscillation frequency with $\beta$ for $k_{n}=1.8$(black dashed), 1(solid black). Here black dotted curve gives $\nu$ in absence of optical lattices. Right panel: Variation of $\delta\nu$ with $\beta$ for $k_{n}=1.5$(solid black), 2(black dotted). In both the panels, we take $G_{-}=1.05$, $x_{0}=0.25$.
The effect of optical lattices on the variation of population imbalance with time is shown in Fig.3. We see that the $Z(t)$ oscillates with time for different $\varphi(0)$ similar to that observed in absence of OLs. However, the effect of the optical lattice is found to change the frequency of oscillation ($\nu$). Specifically, $\nu$ decreases with the increase of $k_{n}$. Like the case of no optical lattice, the value of $\nu$ here also decreases with the increase of $\beta$. The variation of $\nu$ with different lattice and SOC parameters can approximately be calculated from the relation given below.
$\nu=[A\cos 2\varphi(0)(\Delta+\Lambda+B+A\cos^{2}\varphi(0))]^{\frac{1}{2}}.$ (15)
In writing Eq.(15) we have used $\ddot{Z}+\nu^{2}Z=0$. In order to illustrate the effects of optical lattices (OLs) and helicoidal gauge potential on the oscillation of population imbalance in the case of zero Zeeman splitting ($\Delta=0$), we calculate $\nu$ for different values of $\beta$. The results displayed in Fig.4 show that the frequency of JT oscillation decreases as $\beta$ increases for both presence and absence of optical lattices (left panel). Looking closely into the figure we see that the value of $\nu$ for a particular value of $\beta$ also depends on the lattice wavenumber $k_{n}$. More specifically, the value of $\nu$ is larger for a smaller value of $k_{n}$. We note that the deviation in the value of $\nu$ arising due to the variation of $k_{n}$ is appreciable only for smaller values of $\beta$.
With a view to demonstrate how the JT oscillation frequency in presence of optical lattices ($\nu_{2}$) deviates from that in absence of OLs ($\nu_{1}$) we calculate $\delta\nu=\nu_{2}-\nu_{1}$ for different values of $\beta$ ( Fig.4, right panel). We see that $\delta\nu$ gets saturated at the lower value of $\beta$ and it gradually decreases as $\beta$ increases. The saturation of $\delta\nu$ depends on the wavenumber $k_{n}$ of the optical lattice. Specifically, saturated value of $\delta\nu$ is smaller for a larger value of $k_{n}$.
Figure 5: Density evolution of the first component $\rho_{1}=|\psi_{1}|^{2}$ (left panel), second component $\rho_{2}=|\psi_{2}|^{2}$(middle panel) and total $\rho$= $\rho_{1}+\rho_{2}$(right panel) for $\alpha_{s}=4$, $x_{0}=0.25$ and $\Delta=0$.
For zero Zeeman splitting, time evolutions of density profiles of the coupled system are shown in Fig.5. It shows that the density of atomic population is changing between the two components (left and middle panel) periodically with time and the change of population imbalance is symmetric. This implies that maximum population in one component corresponds to minimum population of other component such that the total population remains constant with time (right panel). Note that, for $\Delta=0$, the minimum value of density in both the components is same.
IV Population oscillation for non-zero zeeman splitting
It is seen that the non-zero Zeeman splitting introduces a gap between upper and lower branches of the energy dispersion curve and also causes two non-degenerate energy minima in the lower branch. Thus it is quite interesting to check how the non-zero Zeeman splitting affects the oscillation of population imbalance (Z) in presence and absence of optical lattices. In order of this, we plot the variation of $Z$ with $t$ in Fig.6 for $\Delta=1$. We see that population imbalance oscillates periodically with time and the amplitude of oscillation increases with the increase of $\varphi(0)$ (left panel). However, the oscillation does not occur about $Z=0$ and thus an asymmetry in the maximum population imbalance arises between the two states for any value of $\varphi(0)$. Due to the asymmetry, average value of population imbalance over one period becomes non-zero. This causes the so-called quantum mechanical self-trapping. If we increase $\beta$ keeping $\varphi(0)$ at a fixed value, we see that frequency of Z oscillation reduces and also the asymmetry in the population imbalance increases (right panel). Thus we say that quantum mechanical self-trapping is influenced by helicoidal gauge potential. In presence of optical lattices, we observe similar type of oscillation with slightly different frequency(Fig.7).
Figure 6: Left panel: Variation of population imbalance with $t$ in absence of optical lattices for $\beta=1.6$. Here, solid black and dashed black are drawn for the initial phase differences $\varphi(0)=\pi/8$ and $\pi/16$ respectively. Right panel: It gives $Z(t)$ versus $t$ for $\beta=1.6$(solid black) and $\beta=2$ (black dashed) with $\varphi(0)=\pi/8$. In both the panels we take $\alpha_{s}=4$, $x_{0}=0.25$ and $\Delta=2$.
Figure 7: Left panel: Variation of population imbalance with $t$ in presence of optical lattices for $\beta=1.6$. Here, solid black and dashed black are drawn for the initial phase difference $\varphi(0)=\pi/8$ and $\pi/16$ respectively. Right panel: It gives $Z(t)$ versus $t$ for $\beta=1.6$(solid black) and $\beta=2$ (black dashed) with $\varphi(0)=\pi/8$. In both the panels we take, $\alpha_{s}=4$, $x_{0}=0.25$ and $\Delta=2$.

### v3

@@ Abstract
We consider matter-wave bright solitons in helicoidal spin-orbit coupled Bose-Einstein condensates in optical lattices with a view to study Josephson-type oscillation and self-trapping of population imbalance between two pseudo-spin states. For fixed values of nonlinear interaction and lattice parameters, the population imbalance executes Josephson-type oscillation and the frequency of oscillation decreases with the increase of helicoidal gauge potential. The frequency of oscillation, however, increases in presence of nonlinear optical lattices. We find that the population imbalance can oscillate about a non-zero value and result self-trapping for proper choices of parameters of the system. The self-trapping can be enhanced by the increase of helicoidal gauge potential and by properly tuning the lattice parameters.
@@ I Introduction
Josephson physics describes the tunnelling phenomena that arise due to coupling between two macroscopically phase coherent quantum states such as superconductors separated by thin insulating layer and, weakly coupled Bose-Einstein condensates (BEC) separated by tunnelling layer. The interplay between tunnelling and interaction results supercurrent due to the oscillation of population difference between these two states [1, 2]. This is the so-called Josephson oscillation. A transition from Josephson oscillation to self-trapping occurs if interaction dominates over the tunnelling and causes the population imbalance to oscillate about a non-zero value. This tunnelling oscillation can be explored in building quantum mechanical circuit to design qubit [3], SQUIDS [4] and to engineer quantum states [5].
@@
Two Bose-Einstein condensates in a trap consisting of atoms differing in hyperfine states provide a platform to realize bosonic Josephson junction of a weakly interacting system. A spin-orbit coupled (SOC) atomic BEC created by artificial gauge potentials offers a new possibility to study Josephson oscillation in pseudo-spin 1/2 systems. This type of oscillation in a weakly interacting BEC is a nonlinear generalization of the typical a.c. and d.c. Josephson oscillations. Thus it has received special attention and resulted several interesting experimental [6, 7, 8] and theoretical [9, 10, 11, 12, 13, 14, 15, 16, 17] investigations. Particularly, theoretical investigations in this direction include the studies on Josephson dynamics in two species BECs in double-well potential [10], internal Josephson oscillation between coupled solitons with time varying Raman frequency [17], effects of SOC on Josephson tunnelling between two annular condensates [18], Josephson oscillation of 2D chiral solitons [19] and Josephson-type oscillation in presence of optical lattices [20].
@@
Besides the above mentioned Josephson effects in real space, it has also been predicted theoretically [21] and, realised experimentally [22] momentum-space Josephson effect in SOC-BEC. Studies in this direction are not only limited to atomic-bosonic systems but also extended from exciton-polariton systems [23] to fermionic superfluid gases [24, 25].
@@
Various techniques for generating artificial gauge potentials allow to tune strength and shape of spin-orbit coupling in BECs [26, 27, 28]. For the choice of gauge potential, $A(x)=\sigma n(x)$, where $n(x)=(\cos(2\beta x),\sin(2\beta x),0)$ and $\sigma=(\sigma_{1},\sigma_{2},\sigma_{3})$, one can induce helicoidal shaped non-uniform spin-orbit coupling in BECs. This non-uniform SOC is highly tunable and can increase the effects of intrinsic nonlinearity [26, 27, 28, 29, 30, 31]. A Bose-Einstein condensate with helicoidal SOC features many properties of integrable system such as stable propagation of bright solitons [32], Peregrine solitons [33], bright-dark Peregrine solitons [34], stripe solitons [35], multi-hump solitons [36] and kink solitons [37]. One of the important aspects of the soliton in helicoidal SOC-BECs is that the soliton with any peak amplitude is non-radiative and thus can propagate a large distance [32]. In a recent study it is shown that the helicoidal gauge potential can be used to manipulate modulational instability in BECs [38].
@@
Our objective in this paper is to study Josephson-type (JT) oscillation and self-trapping of atomic population between two solitons in helicoidal spin-orbit coupled quasi-one-dimensional (Q1D) BECs with optical lattices. Particularly, we see that the atomic imbalance between the two solitons oscillates periodically and the frequency of oscillation depends sensitively on the strength of helicoidal gauge potential. The frequency decreases with the increase of helicoidal gauge potential. In presence of optical lattices, the frequency of oscillation further increases [20]. We see that there exists a critical point for the observation of self-trapping. Interestingly, the critical point gets shifted in the parameter space with the change of helicoidal gauge potential and Zeeman-splitting. We note that there is no Rabi coupling in helicoidal SOC-BEC. Therefore, the oscillation of population imbalance occurs due to the helicoidal SOC. In the case of a traditional SOC-BEC, the states are interacting through Rabi coupling [39, 40, 41].
@@
In section II, we present a theoretical model to describe helicoidal SOC-BEC in optical lattices. Based on a variational approach, we derive equations for the center of mass, population imbalance and phase difference of the coupled condensates. In section III, we study the dynamics of population imbalance for a particular location of center of mass in the linar optical lattice. Specifically, we study the effects of nonlinear optical lattices on JT oscillations for different strengths of the helicoidal gauge potentials. In section IV, we find a condition for the observation of quantum mechanical self-trapping and its dependence on the helicoidal gauge potential and nonlinear optical lattice. Finally, we devote section V to summarize the results of our investigations.
@@ II General Formalism
A Bose-Einstein condensate can be treated theoretically by a single mean-field Gross-Pitaevskii equation. In presence of helicoidal spin-orbit coupling, the BEC describes a pseudo-spin 1/2 system and thus it becomes a binary BEC. This system in quasi-one-dimensional (Q1D) geometry is modelled by the following coupled Gross-Pitaevskii equations (GPE) [32, 38].
@@
With a view to consider a helicoidal spin-orbit coupled BEC trapped in linear optical lattices (LOL), we take $\omega_{\perp}\gg\omega_{x}$ such that the effects of harmonic trap is negligible. In addition, we introduce periodicity in the system by modulating the nonlinear mean-field interaction periodically, called nonlinear optical lattice (NOL). The NOLs are given by
@@
where $k_{n}$ is the wavenumber of the NOLs with their strengths $\gamma_{1}$ and $\beta_{1}$ respectively. The NOLs are found very effective in controlling the dynamics and stability of matter-wave solitons in BECs [42, 43, 44, 45, 46, 47]. In order to proceed with the variational approach, we recast the Eqs. (1) and (2) in the following Lagrangian density.
@@
-- $\frac{1}{2}\bigg{|}\frac{d\psi_{1}}{dx}\bigg{|}^{2}-\frac{1}{2}\bigg{|}\frac{d\psi_{2}}{dx}\bigg{|}^{2}-\frac{\Delta}{2}|\psi_{1}|^{2}+\frac{\Delta}{2}|\psi_{2}|^{2}+\frac{\gamma_{0}}{2}|\psi_{1}|^{4}$
$+$ $\frac{{\gamma_{0}}}{2}|\psi_{2}|^{4}+{\beta_{0}}|\psi_{1}|^{2}|\psi_{2}|^{2}-V_{0}\cos(2k_{l}x)|\psi_{1}|^{2}$
-- $V_{0}\cos(2k_{l}x)|\psi_{2}|^{2}+\frac{\gamma_{1}}{2}\cos(2k_{n}x)|\psi_{1}|^{4}$
$+$ $\frac{\gamma_{1}}{2}\cos(2k_{n}x)|\psi_{2}|^{4}+\beta_{1}\cos(2k_{n}x)|\psi_{2}|^{2}|\psi_{1}|^{2}$
@@
A helicoidal spin-orbit coupled BEC supports freely moving bright solitons both in presence and absence of Zeeman splitting [32]. Thus, we adopt the following Gaussian-shaped trial solutions [17].
@@
$+$ $\frac{1}{2\sqrt{2\pi}w}(\gamma_{0}N_{1}^{2}+\gamma_{0}N_{2}^{2}+2\beta_{0}N_{1}N_{2})$
$+$ $\frac{e^{\frac{-k_{n}^{2}w^{2}}{2}}}{2\sqrt{2\pi}w}(\gamma_{1}N_{1}^{2}+\gamma_{1}N_{2}^{2}+2\beta_{1}N_{1}N_{2})\cos(2k_{n}x_{0})$
@@
-- $\frac{Nk_{n}\sin(2k_{n}x_{0})}{2\sqrt{2\pi}w}e^{\frac{-k_{n}^{2}w^{2}}{2}}[G_{+}+G_{-}Z^{2}]$
@@
$\frac{d\varphi}{dt}$ $=$ $2\beta\dot{x}_{0}+\Lambda Z+\frac{NZ}{\sqrt{2\pi}w}e^{\frac{-k_{n}^{2}\omega^{2}}{2}}\cos(2k_{n}x_{0})G_{-}$ (11)
@@
with $\Lambda=N(\gamma_{0}-\beta_{0})/(\sqrt{2\pi})w$, $k_{\pm}=(k_{1}\pm k_{2})/2$, $G_{-}=\gamma_{1}-\beta_{1}$ and $G_{+}=\gamma_{1}+\beta_{1}$. Here, $\varphi=\phi_{1}-\phi_{2}$ and $Z=(N_{1}-N_{2})/N$ represent respectively phase difference and population imbalance between the two solitons. For a weak helicoidal gauge potential we take $k_{-}\approx\beta$. The coupled Eqs. (8)-(11) show that the rate of change of population imbalance depends directly on the initial phase difference and SOC parameters. Clearly, its dependence on the lattice and other parameters comes through $\varphi$. Therefore, it will be quite interesting to check how the population imbalance are affected by the different parameters of the system.
III Helicoidal gauge potentials and JT oscillation
@@
Let us consider that the coupled solitons are located at a particular position in the linear optical lattice (LOL). For these stationary solitons, the center of mass of the coupled system does not move in time and thus $\dot{x_{0}}=0$. In this case, the coupled Eqs. (8)-(11) can be reduced as follows.
$\frac{dZ}{dt}$ $=$ $-A{\sqrt{1-Z^{2}}}\sin(\varphi),$ (12)
$\frac{d\varphi}{dt}$ $=$ $\Lambda Z+\frac{AZ}{\sqrt{1-Z^{2}}}\cos(\varphi)+BZ-\Delta.$ (13)
Here $A=2k_{0}\alpha_{s}e^{-\beta^{2}w^{2}}$ and $B=\frac{NG_{-}e^{\frac{-k_{n}^{2}w^{2}}{2}}}{\sqrt{2\pi}w}\cos(2k_{n}x_{0})$. Note that Eqs. (12) and (13) are obtained by considering weak spin-orbit coupling such that the first term in Eq. (9) is negligible and by choosing the parameters of both the linear and nonlinear lattices such that
$G_{-}\geq\frac{4}{Z_{0}^{2}}\bigg{[}\sqrt{2\pi}w\frac{V_{0}k_{l}\sin(2k_{l}x_{0})}{\sin(2k_{n}x_{0})}e^{-\frac{w^{2}}{2}(2k_{l}^{2}+k_{n}^{2})}-\frac{G_{+}}{4}\bigg{]}$ (14)
with $10<Z_{0}<$ 1. Under this condition $k_{+}$ remains approximately constant with time. Therefore, Eqs. (12) and (13) can describe the variation of population imbalance with a change of relative phase between the two spin states in presence of both linear and nonlinear optical lattices (NOL).
@@
With a view to study population imbalance dynamics, we consider that the components are miscible ($\gamma_{0}>\beta_{0}$) and solve Eqs. (12) and (13) numerically for different values of initial phase differences and display the results in Fig. 1. We see that, in absence of NOLs the population imbalance oscillates with time and the amplitude of oscillation depends sensitively on the initial value of $\varphi$ (left panel). More specifically, the amplitude increases with the increase of $\varphi(0)$ for a fixed value of $\beta$. This type of oscillation is often termed as Josephson-type (JT) oscillation [16, 20, 17]. The frequency of oscillation decreases with the increase of $\beta$ (right panel) and thus we say that population imbalance dynamics can be decelerated with the help of helicoidal gauge potential.
@@
Figure 1: Left panel: Variation of population imbalance with $t$ in absence of nonlinear optical lattices for $\beta=1.6$. Here, dashed blue and dotted black are drawn for the initial phase differences $\varphi(0)=\pi/8$ and $\pi/16$ respectively. Right panel: It gives $Z(t)$ versus $t$ for $\beta=1.6$ (solid blue) and $\beta=2$ (dashed black) with $\varphi(0)=\pi/8$. In both the panels, we take $\alpha_{s}=4$, $x_{0}=0.25$, $\Delta=0$, $B=0$, $k_{0}=1$ and $\Lambda=0.04$.
@@
Figure 2: Left panel: Variation of population imbalance with $t$ in presence of nonlinear optical lattices for $\beta=1.6$. Here, solid blue and dashed black are drawn for the initial phase differences $\varphi(0)=\pi/8$ and $\varphi(0)=\pi/16$ respectively. Right panel: It gives $Z(t)$ versus $t$ for $\beta=1.6$ (solid blue) and $\beta=2$ (solid black) with $\varphi=\pi/8$. In both the panels, we take $\alpha_{s}=4$, $x_{0}=0.25$, $G_{-}=1.05$, $V_{0}=2$, $k_{l}=5$, $k_{n}=1.8$, $\Delta=0$, $k_{0}=1$ and $\Lambda=0.04$.